OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = 9x + 3 and g(x) = Quantity x minus three divided by nine

OpenStudy (anonymous):

f(x) = 9x + 3 and g(x) = (x -3)/9 So just plug one in for x into the other. f(g(x)) = 9((x-3)/9)+3 g(f(x)) = ((9x+3)-3)/9

OpenStudy (anonymous):

thanks